OpenStudy (anonymous):

Can you please check my work? Two particles start out at t = 0 at he origin. One moves along the positive x-axis at the rate of sin(3pi*t) meters per second and the other moves along the positive y-axis at he rate of 1/(1 + t^2) meters per second. What is the net distance between them after 10 seconds?

OpenStudy (anonymous):

I found the integral of sin(3pi*t) to be (-cos(3pi*t))/3pi

OpenStudy (anonymous):

I then evaluated the equation @ t=10 and got -1

OpenStudy (anonymous):

Then I did the same for the second equation and got 1.47113

OpenStudy (anonymous):

Then I used the distance formula and got 1.7788 as my answer. It this right?

OpenStudy (anonymous):

Very close, but not quite! \[ x' = sin(3\pi t) \]\[x=-\frac{cos(3\pi t)}{3\pi} + c\] Now we apply our initial condition: x(0)=0: \[x(0)=0=-\frac{1}{3\pi} + c\]\[c=\frac{1}{3\pi}\]Combining this, we see: \[x=\frac{1-cos(3\pi t)}{3\pi}\] and \[x(10)=\frac{1-cos(30\pi)}{3\pi} = \frac{0}{3\pi}=0\]This may seem odd, but as -1 <= sin(x) <= 1, the speed can be negative implying it's going backwards. Now for y: \[y'=\frac{1}{1+t^2}\]\[y=tan^{-1}(t) + c \]Applying the condition y(0)=0, we see c=0 so \[y=tan^{-1}(t)\] So \[y(10)=tan^{-1}(10) = 1.471127674\]. As x=0, our value for y(10) is the distance :)

OpenStudy (anonymous):

I see what I did wrong. I neglected the constant. Thanks for your help.

OpenStudy (anonymous):

Yeah it's an easy mistake to make, you were so close. No problem!